OpenStudy (lilpeter504):

Can someone help me with alge 2 substitution and elimination? : big bracket { x + 2y = -7 { 2x - 3y = 0

OpenStudy (anonymous):

yes i can

OpenStudy (lilpeter504):

Good, thank you! :)

OpenStudy (anonymous):

x + 2y = -7 (equation 1) 2x - 3y = 0 (equation 2) first for equation 1 x + 2y = -7 x = -7 - 2y (equation 3) substitute this equation 3 to equation 2, then 2x - 3y = 0 2(-7 - 2y ) - 3y =0 -14 - 4y -3y = 0 -14 -7y = 0 -7y = 14 y = 14/(-7) y = -2

OpenStudy (anonymous):

from equation 1 x + 2y = -7 x + 2 (-2) = -7 x -4 = -7 x = -7 +4 x = -3

OpenStudy (lilpeter504):

Wait, why did you separate the x and not the y, i should be getting the y alone correct?

OpenStudy (lilpeter504):

If I were to separate the y, the answer is y = 1/2x-3.5 for equation 1 correct?

OpenStudy (lilpeter504):

I got a different answer when i did substitution on the way I did it

OpenStudy (lilpeter504):

(x,y) = (0,-3.5) I am not entirely sure if its correct